# DuoSpaceNet: Leveraging Both Bird’s-Eye-View and Perspective View Representations for 3D Object Detection

### v1

@@ Abstract
Recent advances in multi-view camera-only 3D object detection either rely on an accurate reconstruction of bird’s-eye-view (BEV) 3D features or on traditional 2D perspective view (PV) image features. While both have their own pros and cons, few have found a way to stitch them together in order to benefit from “the best of both worlds”. To this end, we explore a duo space (i.e., BEV and PV) 3D perception framework, in conjunction with some useful duo space fusion strategies that allow effective aggregation of the two feature representations. To the best of our knowledge, our proposed method, DuoSpaceNet, is the first to leverage two distinct feature spaces and achieves the state-of-the-art 3D object detection and BEV map segmentation results on nuScenes dataset.
Figure 1: Comparison of different image-only 3D perception frameworks. (a) Top: Bird’s-eye-view-based (BEV-based) methods (b) Middle: Perspective-view-based (PV-based) 3D detection-only methods (c) Bottom: DuoSpaceNet (ours) where 3D detection benefits from both 3D BEV and 2D PV feature space. Dense segmentation is also doable.
Figure 2: Example of differences between bird’s-eye-view (BEV) features and perspective view (PV) features. (a) Left: The original front-facing RGB image. (b) Middle: PV heat map. (c) Right: BEV heat map. Both heat maps are produced by our final model. Different hues and brightness represent different responses. Black represents low or no response. We find that each feature space has its own strengths. In this example, it is easier to reason object positions (e.g., the relative positions of the two leading vehicles in red dashed boxes) by looking into the BEV heat map. Meanwhile, none of the vehicle responses in BEV space overlap. When it comes to the PV heat map, it seems that more detailed semantics are preserved with a higher resolution, making it more desired for object attribute prediction. Also, it is free from camera ray-like artifacts, thus being less noisy.
@@ 1 Introduction
3D detection and segmentation via multi-view images undergoes active research [35, 15, 22] due to its usefulness for applications such as autonomous driving. While there are many LiDAR-based 3D detection and segmentation methods which yield better results, camera-only methods still have their unique advantages. For example, camera-only perception systems are generally cost-friendly to deploy and have a higher resolution for distant objects.
@@
Most existing methods [35, 15, 22, 18, 5] fall into one of the following two categories: bird’s-eye-view-based (BEV-based) methods or perspective-view-based (PV-based) methods, as depicted in Fig. 1. BEV-based methods [15, 5] spend great efforts on generating a bird’s-eye-view (BEV) feature map via a 2D to 3D lifting strategy, usually based on camera projection or unprojection model. Each element in the BEV feature map corresponds to the feature of a mapped 3D location. Subsequent detection and segmentation tasks are performed directly in BEV feature space. PV-based 3D detection methods [35, 18], however, directly build their detection heads on top of 2D image features. These methods generally utilize a handful of sparse detection proposals, often in the hundreds. For 3D detection, visual cues are directly gathered from sparse locations on PV image features where the proposals are anchored at. PV-based methods are specialized for object detection. Generally speaking, they are more computationally efficient.
While both BEV-based methods and PV-based methods seem to work well, some of their shortcomings are outstanding. When it comes to BEV-based methods, since multi-view pose parameters guide the feature lifting process, BEV features tend to have a better understanding of the underlying 3D geometry of a scene, which is crucial for estimating the 3D position of an object. BEV features are more robust to object scale variations and are free from overlap issues caused by perspective projection. However, subtle visual cues may be compromised due to coarse grid granularity, downsampling or interpolation. Additionally, because of the fixed-size nature of the BEV feature map, objects beyond the pre-set BEV range are destined to be disregarded, even if they possess a very large speed and will interact with the autonomous vehicle shortly. In contrast, PV-based methods seem to maximize the efficiency for 3D object detection by attending to 2D PV features sparsely and are capable of detecting objects at any distance theoretically. However, they lack certain functionalities and extensibilities such as handling map segmentation as well as multi-modal inputs. Moreover, they also tend to exhibit inferior 3D perception performance compared to BEV-based methods, as 3D information can only be implicitly inferred from PV features.
Take Fig. 2 as an instance, it is apparently more difficult to retrieve 3D spatial clues from the PV than the BEV feature map. Moreover, being perspective, the PV feature map suffers from occlusion and perspective distortion, whereas we can have an unobstructed and rectified view of objects in BEV space. However, BEV feature maps lack fine-grained details, especially for small objects such as pedestrian and cyclist, due to resolution, and may contain camera ray-like artifacts from back-projection, whereas PV features are free from these issues.
Based on the above reasonings, it looks like BEV-based and PV-based methods are rather complementary. In order to bridge the gap between current BEV-based and PV-based frameworks to preserve “the best of both worlds”, in this paper, we propose DuoSpaceNet, a new paradigm that jointly tackles 3D object detection and map segmentation tasks via both BEV and PV feature representations. In our network, both PV features and BEV features are retained and fed into our Duo Space Decoder for object detection, a variant of the transformer decoder used in Deformable DETR [40]. To maintain the uniqueness of each feature space while creating a unified representation for individual 3D object, each object query comprises duo space content embeddings from both PV and BEV spaces, alongside a shared pose embedding that represents its real-world 3D pose. The decoder incorporates space-specific cross-attention layers to refine the duo space object queries using features from both spaces. To maximize the distinctiveness between BEV and PV features, feature divergence enhancement is introduced as the finishing touch of our BEV feature generation process. A temporal version of DuoSpaceNet is also established to show our framework’s adaptability from single to multiple frames. To achieve this, we propose a unique duo space temporal solution which elegantly processes temporal inputs via duo space temporal queries. For map segmentation, we append a convolution-based segmentation heads after BEV feature generation, and predict each map category separately.
In short, our novelties and contributions are as follows.
  • To the best of our knowledge, we are the first to identify and address the complementary nature of feature modeling in bird’s-eye-view (BEV) space and perspective view (PV) space. Our proposed model, namely DuoSpaceNet, is the first approach to perform 3D detection based on a unique multi-feature-space paradigm.
  • To accomplish our duo space framework, we propose: 1) a duo space decoder for duo space feature extraction and fusion; 2) a divergence feature enhancement design to increase inter-space distinctiveness; and 3) a duo space temporal modeling method to extend our paradigm to a multi-frame setting.
  • On nuScenes dataset [1], we conduct extensive experiments and ablation studies to verify the performance and effectiveness of our proposed ideas. We are able to demonstrate that DuoSpaceNet achieves the state-of-the-art results on both 3D object detection and map segmentation benchmarks, as well as the necessity of each individual design that we propose.
@@ 2.1 BEV-Based Multi-View 3D Perception
Tackling multi-view 3D object detection task by bird’s-eye-view (BEV) representations has been a popular trend in autonomous driving industry. Following LSS [30], BEVDet [10, 9] and BEVDepth [14] unproject 2D image features into an explicit BEV feature map using dense depth predictions. M2BEV [36] and SimpleBEV [6] improve the efficiency of 2D to 3D BEV transformation by assuming a uniform distribution of depths when doing camera back-projection. Surprisingly, BEV-based methods can work without the non-trivial dense depth estimation. BEVFormer series [15, 37] model dense BEV feature map with per-dataset queries optimized via deformable attention [40]. Going on, most recent BEV-based 3D detection methods [28, 29, 5, 41] shift their endeavors to improve temporal designs rather than fundamentals of BEV feature representations.
Another focus of BEV-based 3D perception is to handle the map segmentation task. Early works [25, 30, 27] tend to treat it as an individual task, while recent works, such as M2BEV and BEVFormer, explore the potential of jointly tackling object detection and map segmentation tasks by multi-task learning.
@@ 2.2 PV-Based Multi-View 3D Perception
Starting with DETR3D [35], the landscape of multi-view perspective-view-based (PV-based) 3D object detection leans towards sparse query refinement with set-to-set matching loss. PETR [21, 22] improves transformer-based detection decoders with 3D position-aware image features. Recently, Sparse4D [18] further extends this track with the introduction of 4D anchors, allowing intuitive ego-motion and object motion compensation. We also find it useful and derive our query design from it. Similar to the BEV-based 3D detection track, current state-of-the-art PV-based methods, such as [19, 34, 20], also focus on improving temporal modeling techniques by temporal query propagation.
@@ 3.1 Duo Space Features
Our model benefits from the contrastiveness between the two feature representations. Since our lifting method is parameter-free, its functionality can be viewed as rearranging PV features given priors (e.g., camera poses) on the 3D geometry of a scene. Therefore, it has minimal effects on diverging the feature contents. To increase the heterogeneity of BEV features w.r.t. PV features, we propose a simple yet effective divergence enhancement stage. It consists of three 3D convolution layers (Conv3Ds) and three 2D convolution layers (Conv2Ds). First, we apply Conv3Ds on $F_{voxel}$ to improve 3D geometry awareness in a learning-based fashion. After $F_{voxel}$ is flattened along its $Z$ dimension, Conv2Ds are applied for further BEV-level refinement, yielding final $F_{BEV}$.
@@ 3.2 Duo Space Decoder
Suppose we have $k$ object queries, $\{Q^{i}\}_{i=1}^{k}$. Each consists of a pose embedding, $Q^{i}_{Pose}$, and duo space content embedding for both BEV and PV spaces, $Q^{i}_{BEV}$ and $Q^{i}_{PV}$, respectively. Adapted from [18], each $Q^{i}_{Pose}$ is encoded from a 3D pose vector $\mathcal{P}_{i}$, which contains attributes with physical meanings, including $x,y,z$ in the vehicle coordinate system, the width, length, height, orientation and the velocity of the object the query is associated with. While $Q^{i}_{BEV}$ and $Q^{i}_{PV}$ contain high-level content features in BEV space and PV space respectively. In each layer of the duo space decoder, first, a pose encoder consisting of several FC layers is used to encode $\{\mathcal{P}_{i}\}_{i=1}^{k}$ into high dimensional latent representations, dubbed $\{\texttt{Enc}\left(\mathcal{P}_{i}\right)\}_{i=1}^{k}$, which will serve as learnable positional encodings in the subsequent attention layers. To unify the 3D pose of each object query across BEV and PV spaces, we generate a shared pose embedding,
@@
For multi-head cross-attention layers $\texttt{MHCA}_{BEV}(...)$ and $\texttt{MHCA}_{PV}(...)$, each of them will only act on their corresponding features using corresponding inputs. In BEV space, it can be represented as
@@ 3.3 Duo Space Temporal Modeling
BEV-based 3D detection methods [9, 15] typically utilize temporal inputs by stacking temporal BEV feature maps. Offsets are determined either with motion compensation or in a learnable manner (e.g., deformable attention) or both combined. Meanwhile, PV-based methods [18, 22] generally infuse temporal information into object queries. Therefore, the difference between BEV-based and PV-based temporal methods brings challenges to temporal design in our duo space paradigm. In this section, we present a unified temporal solution for both spaces via temporal duo space queries, illustrated in Fig. 4. Concretely, for a fixed temporal length $l$, each object is represented by $l$ temporal duo space queries, each comprising a temporal pose of the underlying object and a shared content embedding. Temporal poses are deduced by applying both ego- and object-motion compensation on the object’s current pose vector. The queries are fed into the space-specific cross-attention layer corresponding to their space. Each query only attends to features from a specific timestamp associated with its temporal pose. Subsequently, results produced by $l$ temporal queries are recurrently aggregated via multi-layer perceptron (MLP) into a single refined prediction per object. Our solution elegantly operates symmetrically across BEV and PV spaces. Please refer to the supplementary materials for more details.
@@ 4 Experiments
Implementation details. For both single-frame and multi-frame 3D detection experiments, if not specified otherwise, we completely follow the hyperparameter settings in [18], including the learning rate and its schedule, data augmentation, loss functions and anchor initialization. For full model experiments, the BEV feature map is sized $200\times 200$, the number of duo space queries is $900$ and the number of decoder layers is 4. All layers have identical settings with $8$ attention heads in both self-attention and cross attention layers. For deformable cross attention layers, we compute $16$ offsets per query. For multi-frame experiments, we use 4 adjacent frames (including the current frame) as temporal input. For all ablation studies, we use ResNet-50 [7], $100\times 100$ BEV feature map (if applicable), $800\times 320$ input images and a 2-layer decoder, trained for 12 epochs. For map segmentation, we follow PETRv2 [22] to transform map layers from the global reference frame into the ego frame, and generate two $200\times 200$ ground truth segmentation masks for $drivable\ area$ and $lane\ boundary$ respectively.
@@
Table 2: Map segmentation on the nuScenes val set.
@@
Our 3D detection results on nuScenes val set are shown in Tab. 1. Compared with other state-of-the-art single-/multi-frame methods, our method consistently outperforms others on mAP. Specifically, we achieve 1.7% mAP gain over the state-of-the-art PV-based method Sparse4D [18] and 2.4% mAP gain over the state-of-the-art BEV-based method BEVFormer-S [15], using the single-frame setup. The same is true for multi-frame results. Among all methods, DuoSpaceNet achieves the lowest mATE by a large margin, suggesting that our duo space design helps the model understand 3D scenes better. When it comes to other metrics, although our method does not achieve 1st place for some entries, we argue that on average our model surpasses others based on the NDS measurement. We also report our results on nuScenes test set in Tab. 3. Compared with PolarFormer-T [11], DuoSpaceNet achieves a considerable 1.2% mAP gain and 2.6% NDS gain. Note that different methods use different training strategy on the test set (e.g., longer training schedules, more temporal frames, etc.). Nonetheless, our model is capable of achieving competitive results against other state-of-the-art models.

### v3

@@ Abstract
Multi-view camera-only 3D object detection largely follows two primary paradigms: exploiting bird’s-eye-view (BEV) representations or focusing on perspective-view (PV) features, each with distinct advantages. Although several recent approaches explore combining BEV and PV, many rely on partial fusion or maintain separate detection heads. In this paper, we propose DuoSpaceNet, a novel framework that fully unifies BEV and PV feature spaces within a single detection pipeline for comprehensive 3D perception. Our design includes a decoder to integrate BEV–PV features into unified detection queries, as well as a feature enhancement strategy that enriches different feature representations. In addition, DuoSpaceNet can be extended to handle multi-frame inputs, enabling more robust temporal analysis. Extensive experiments on the nuScenes dataset show that DuoSpaceNet surpasses both BEV-based baselines (e.g., BEVFormer) and PV-based baselines (e.g., Sparse4D) in 3D object detection and BEV map segmentation, verifying the effectiveness of our proposed design.
Equal contribution. Corresponding author. This work was done at TuSimple (now CreateAI).
Figure 1: Comparison of different image-only 3D perception frameworks. (a) Top: Bird’s-eye-view-based (BEV-based) methods. (b) Middle: Perspective-view-based (PV-based) 3D detection-only methods. (c) Bottom: DuoSpaceNet (ours) where 3D detection benefits from both 3D BEV and 2D PV feature space.
Figure 2: Comparison of bird’s-eye-view (BEV) and perspective-view (PV) features. (a) Left: The original front-facing RGB image. (b) Middle: PV heatmap. (c) Right: BEV heatmap. Both heatmaps are produced by our final model. Different hues and brightness levels represent various response intensities with black indicating minimal or no response. We highlight two car instances, each shown with its original image patch as well as corresponding PV and BEV heatmaps. The BEV heatmap in (c) makes it easier to interpret 3D positions—for instance, the relative positions of leading vehicles—without overlap issues. In contrast, the PV heat map in (b) preserves finer semantic details at higher resolution, which benefits attribute prediction.
@@ 1 Introduction
3D detection using multi-view images represents an active area of research [48, 18, 11, 54, 10, 9, 56, 24, 22, 38, 27, 23, 47, 8], driven by critical applications in autonomous driving [12] where it substantially improves downstream tasks such as object tracking [41, 17] and motion prediction [4, 63, 60, 61]. Although LiDAR-based methods [59, 3, 58] often deliver superior performance for 3D detection, camera-only solutions provide distinct advantages: they are typically more cost-effective to deploy, perform reliably in adverse weather conditions such as rain or snow [62], and offer higher resolution at greater distances with long focal length.
@@
Most existing multi-view camera-only 3D object detection methods [48, 18, 27, 21, 5] fall into two broad categories: bird’s-eye-view-based (BEV-based) and perspective-view-based (PV-based), as illustrated in Fig. 1. BEV-based methods [18, 5] generate a bird’s-eye-view feature map through a 2D-to-3D lifting process—often leveraging camera projection or unprojection—where each element of the BEV map corresponds to a specific 3D location. Conversely, PV-based methods [48, 21] construct 3D detection heads atop 2D image features, typically relying on a limited set of sparse detection proposals. Visual cues are derived from the sparse feature locations where these proposals are anchored. In general, PV-based methods prioritize object detection efficiency and computational speed.
While both BEV-based and PV-based methods can be effective, each exhibits notable limitations. In BEV-based methods, multi-view pose parameters guide the feature-lifting process, enhancing 3D geometric awareness and ensuring robustness to object-scale variations. They also avoid overlap artifacts arising from perspective projection. However, subtle visual cues can be lost due to coarse grid resolutions or interpolation. Moreover, the fixed-size nature of the BEV feature map discards objects outside its predefined coverage, even if they are rapidly approaching the ego-vehicle. By contrast, PV-based methods often maximize efficiency by sparsely attending to 2D PV features, allowing for theoretical detection at any distance. Yet, these methods typically rely on implicitly inferring 3D cues from 2D features, leading to weaker 3D perception. As shown in Fig. 2, retrieving spatial information from PV features is more difficult because of occlusion and perspective distortion, whereas BEV provides an unobstructed, rectified view. Still, BEV maps may omit fine-grained details for small objects (e.g., pedestrians or cyclists) and are subject to camera ray-like artifacts and limited coverage—issues that PV features do not face.
Some methods have attempted partial fusion of BEV and PV representations, yet they often depend on separate detection heads or derive visual cues predominantly from PV features. For example, BEVFormerv2 [57] combines BEV and PV features but maintains distinct detection heads, while the PETR series [26, 28] focuses on integrating BEV positional encodings into PV features. These partial approaches underscore the need for a more holistic solution to fully harness the complementary strengths of BEV and PV.
We introduce DuoSpaceNet, a holistic 3D perception framework that bridges the gap between bird’s-eye-view (BEV) and perspective-view (PV) approaches by leveraging both representations for 3D object detection and map segmentation. The core novelty of DuoSpaceNet is its truly integrated dual-space (also written as duo space) representation, where each object query contains both BEV and PV embeddings. Unlike systems that merely combine features from separate heads, our method preserves the distinctiveness of each representation while merging them in a single, unified decoder for robust 3D perception.
Our architecture retains BEV and PV features and processes them with the Duo Space Decoder—a modified version of the transformer decoder from Deformable DETR. Each object query includes dual content embeddings from both spaces, along with a shared pose embedding representing the object’s real-world 3D position. Within the decoder, space-specific cross-attention layers refine dual-space object queries, enabling synergy between geometric and semantic cues. We further propose a feature divergence enhancement that maximizes the distinctiveness of BEV and PV features, leading to more discriminative representations.
Beyond singe-frame detection, DuoSpaceNet supports a temporal extension that elegantly infuses multi-frame inputs into temporal queries, demonstrating our framework’s adaptability. For map segmentation, we append a segmentation head after BEV feature generation, with separate channels for each map category. By fusing both geometric and semantic insights in a single pipeline, DuoSpaceNet offers a comprehensive solution for multi-task 3D perception, setting it apart from existing methods that rely on partial fusion or separate detection heads.
Our main contributions are as follows:
  • Integrated Duo Space Framework: We present DuoSpaceNet, which unifies BEV and PV representations within a single detection pipeline. This design features a duo space decoder for comprehensive BEV–PV fusion and a divergence feature enhancement technique that preserves the strengths of each view while enriching the combined representation.
  • Temporal Extension: Our method also extends to temporal dimension with motion-compensated temporal queries, assisting 3D detection with multi-frame context.
  • Empirical Validation on nuScenes: Through extensive experiments on the nuScenes [1] dataset, DuoSpaceNet outperforms both PV-only and BEV-only baselines, as well as other strong methods, in 3D detection and segmentation. Ablation studies further confirm the effectiveness of each proposed component.
@@ 2.1 BEV-Based Multi-View 3D Perception
Tackling multi-view 3D object detection using bird’s-eye-view (BEV) representations has become popular in autonomous driving. Methods like LSS [37], BEVDet [10, 9], and BEVDepth [16] project 2D image features into BEV feature maps using dense depth predictions. M2BEV [54] and SimpleBEV [6] enhance efficiency by assuming uniform depth during back-projection. Notably, BEV methods such as BEVFormer series [18, 56, 38] avoid dense depth estimation by modeling BEV features with dataset-specific queries via deformable attention [68]. BEVFormer v2 [56] incorporates an auxiliary 3D detection head, while OCBEV [38] focuses on instance-level temporal fusion. Recent methods [34, 35, 5, 69] primarily improve temporal modeling rather than fundamental BEV representations.
BEV-based methods also address map segmentation tasks. VED [31] uses a variational encoder-decoder to generate BEV features, and VPN [33] employs an MLP-based transformation. Transformer-based approaches [65, 36] utilize cross-attention mechanisms to produce BEV representations. Recent geometric reasoning methods such as LSS [37] explicitly estimate depth distributions, while methods like M2BEV and BEVFormer leverage camera parameters and multi-task learning for enhanced detection and segmentation performance.
@@ 2.2 PV-Based Multi-View 3D Perception
Starting with DETR3D [48], the landscape of multi-view perspective-view-based (PV-based) 3D object detection leans towards sparse query refinement with set-to-set matching loss. Following studies like PETR [26, 27] and CAPE [55] improve transformer-based detection decoders with 3D position-aware image features. Recently, Sparse4D [21] further extends this track with the introduction of 4D anchors, allowing intuitive ego-motion and object motion compensation. We also find it useful and derive our query design from it. Similar to the BEV-based 3D detection track, current state-of-the-art PV-based methods, such as [22, 47, 24], also focus on improving temporal modeling techniques by temporal query propagation.
2.3 Unifying BEV and PV Detection
While several approaches have explored the fusion of Bird’s-Eye-View (BEV) and Perspective-View (PV) features for 3D detection, our proposed method stands apart by employing a fully shared detection head. BEVFormerv2 [57] pioneered the combination of BEV and PV queries to enhance BEV detection performance, but maintains separate detection heads for BEV and PV tasks. Furthermore, their detection queries contain either PV or BEV latent features exclusively, rather than integrating both. The PETR series [26, 28] and StreamPETR [46] represent an alternative approach utilizing 3D queries and 2D PV features to detect 3D objects. However, these methods differ fundamentally from our work as their 3D queries contain only position embeddings, with all visual information derived solely from 3D PV image features. In contrast, our 2D-3D fused queries explicitly generate and incorporate BEV features, obtaining more comprehensive cues from both feature representations. MVF [66] presents perhaps the closest parallel to our approach, proposing a LiDAR-based 3D detection pipeline that leverages features from both 3D BEV space and 2D range image space. However, their method does not naturally extend to query-based camera-only detection systems like ours.
@@ 3.1 Duo Space Features
Our model benefits from the contrastiveness between the two feature representations. Since our lifting method is parameter-free, its functionality can be viewed as rearranging PV features given priors (e.g., camera poses) on the 3D geometry of a scene. Therefore, it has minimal effects on diverging the feature contents. To increase the heterogeneity of BEV features w.r.t. PV features, we propose a simple yet effective divergence enhancement stage.
It consists of three 3D convolution layers (Conv3Ds) and three 2D convolution layers (Conv2Ds). First, we apply Conv3Ds on $F_{voxel}$ to improve 3D geometry awareness in a learning-based fashion. After $F_{voxel}$ is flattened along its $Z$ dimension, Conv2Ds are applied for further BEV-level refinement, yielding final $F_{BEV}$.
@@ 3.2 Duo Space Decoder
Suppose we have $k$ object queries, $\{Q^{i}\}_{i=1}^{k}$. Each consists of a pose embedding, $Q^{i}_{Pose}$, and duo space content embedding for both BEV and PV spaces, $Q^{i}_{BEV}$ and $Q^{i}_{PV}$, respectively.
Adapted from [21], each $Q^{i}_{Pose}$ is encoded from a 3D pose vector $\mathcal{P}_{i}$, which contains attributes with physical meanings, including $x,y,z$ in the vehicle coordinate system, the width, length, height, orientation and the velocity of the object the query is associated with. While $Q^{i}_{BEV}$ and $Q^{i}_{PV}$ contain high-level content features in BEV space and PV space respectively. In each layer of the duo space decoder, first, a pose encoder consisting of several FC layers is used to encode $\{\mathcal{P}_{i}\}_{i=1}^{k}$ into high dimensional latent representations, dubbed $\{\texttt{Enc}\left(\mathcal{P}_{i}\right)\}_{i=1}^{k}$, which will serve as learnable positional encodings in the subsequent attention layers.
To unify the 3D pose of each object query across BEV and PV spaces, we generate a shared pose embedding,
@@
For multi-head cross-attention layers $\texttt{MHCA}_{BEV}(...)$ and $\texttt{MHCA}_{PV}(...)$, each of them will only act on their corresponding features using corresponding inputs.
In BEV space, it can be represented as
@@ 3.3 Duo Space Temporal Modeling
BEV-based 3D detection methods [9, 18] typically utilize temporal inputs by stacking temporal BEV feature maps. Offsets are determined either with motion compensation or in a learnable manner (e.g., deformable attention) or both combined. Meanwhile, PV-based methods [21, 27] generally infuse temporal information into object queries. Therefore, the difference between BEV-based and PV-based temporal methods brings challenges to temporal design in our duo space paradigm. In this section, we present a unified temporal solution for both spaces via temporal duo space queries, illustrated in Fig. 4. Concretely, for a fixed temporal length $l$, each object is represented by $l$ temporal duo space queries, each comprising a temporal pose of the underlying object and a shared content embedding. Temporal poses are deduced by applying both ego- and object-motion compensation on the object’s current pose vector.
The queries are fed into the space-specific cross-attention layer corresponding to their space. Each query only attends to features from a specific timestamp associated with its temporal pose. Subsequently, results produced by $l$ temporal queries are recurrently aggregated via multi-layer perceptron (MLP) into a single refined prediction per object. Our solution elegantly operates symmetrically across BEV and PV spaces. Please refer to the supplementary materials for more details.
@@ 4 Experiments
Implementation details. For both single-frame and multi-frame 3D detection experiments, unless specified otherwise, we follow the hyperparameter settings in [21], including learning rate and schedule, data augmentation, loss functions, and anchor initialization. For full model experiments, the BEV feature map is sized $200\times 200$, the number of duo space queries is $900$ and the number of decoder layers is 4. All layers have identical settings with $8$ attention heads in both self-attention and cross attention layers. For deformable cross attention layers, we compute $16$ offsets per query. For multi-frame experiments, we use 4 adjacent frames (including the current frame) as temporal input. For all ablation studies, we use ResNet-50 [7], $100\times 100$ BEV feature map (if applicable), $800\times 320$ input images and a 2-layer decoder, trained for 12 epochs. For map segmentation, we follow PETRv2 [27] to transform map layers from the global reference frame into the ego frame, and generate two $200\times 200$ ground truth segmentation masks for $drivable\ area$ and $lane\ boundary$ respectively.
@@
[…]
BEVDet BEV 69.5M 1498.8G 0.339 0.389
BEVFormer-S BEV 66.6M 1705.5G 0.375 0.448
[…]
DuoSpaceNet BEV+PV 64.8M 1771.7G 0.399 0.462
Table 2: Comparison on model complexity in terms of the number of parameters and the number of floating-point operations.
@@
Table 3: Map segmentation on the nuScenes val set.
Our 3D detection results on nuScenes val set are shown in Tab. 1. Compared with other state-of-the-art single-/multi-frame methods, our method consistently outperforms others on mAP. Specifically, we achieve 1.7% mAP gain over the state-of-the-art PV-based method Sparse4D [21] and 2.4% mAP gain over the state-of-the-art BEV-based method BEVFormer-S [18], using the single-frame setup. The same is true for multi-frame results. Among all methods, DuoSpaceNet achieves the lowest mATE by a large margin, suggesting that our duo space design helps the model understand 3D scenes better. When it comes to other metrics, although our method does not achieve 1st place for some entries, we argue that on average our model surpasses others based on the NDS measurement. We also report our results on nuScenes test set in Tab. 4. Compared with PolarFormer-T [13], DuoSpaceNet achieves a considerable 1.2% mAP gain and 2.6% NDS gain. Note that different methods use different training strategies on the test set (e.g., longer training schedules, more temporal frames, etc.). Nonetheless, our model is capable of achieving competitive results against other state-of-the-art models.
@@
We also compare our model complexity against other state-of-the-art BEV-only or PV-only methods using input images of size $1600\times 640$ under the single-frame setting. For all models, we test them on the same machine using DeepSpeed Flops Profiler [^1] https://www.deepspeed.ai/tutorials/flops-profiler/. As shown in Tab. 2, under similar model sizes, DuoSpaceNet significantly outperforms BEVDet [10] and BEVFormer-S. It is also slightly better than Sparse4D, yet still capable of handling dense segmentation tasks.